OpenStudy (sissyedgar):

Ummmmm...... Use the Distributive Property and simplify the expression. 6 + 4(3x – 2) + 5 A. 12x + 9 B. 3x + 13 C. 12x + 3 D. 7x + 9

pooja195 (pooja195):



OpenStudy (anonymous):

6 + 4(3x-2) + 5 First, multiply 4 to (3x-2) 6 + 4(3x-2) + 5 = 6 + 12x - 8 + 5 Next, group similar terms and compute them = [12x] + [6 - 8 +5] = 12x + 3 (D)

pooja195 (pooja195):

Plagerized answer.

OpenStudy (sissyedgar):

@YK.PAKI 12x + 3 is (C) not (D) lol, its fine tho...

OpenStudy (sissyedgar):

@YK.PAPI *

OpenStudy (anonymous):

itsnot plagerizedi really thought that was the answer

OpenStudy (sissyedgar):

it is the rite answer {i guess}, pooja ment not 2 give the answer at all.

OpenStudy (anonymous):

oh ok sorry i jus seen every one giving answers out and im new to openstudy

OpenStudy (anonymous):

thank you very much @pooja195 @sissyedgar

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)